OpenStudy (anonymous):

what is positive exponents? I can find negative exponents but not positive exponents

OpenStudy (anonymous):

a positive exponent is like the 2 in x^2

OpenStudy (anonymous):

16x^{4}+12x ^{3} +14x ^{2}/2x ^{3} how do i find the positive exponents while performing the division

OpenStudy (amistre64):

subtract the bottom fromthe top; and if its negative throw it back under

OpenStudy (anonymous):

If you get negative exponents you can just move it to the other side of the fraction bar.. \[a^{-3} = \frac{1}{a^3}\]\[\frac{1}{b^{-5}} = b^5\]\[\frac{c^{-2}}{d^{-3}} = \frac{d^3}{c^2}\]